OpenStudy (anonymous):

simplify -(2x-3y)+5x+2-3x+3y+2 A.4 B.4x +4 C.6y +4 D.4x + 6y +4

OpenStudy (anonymous):

–(2x – 3y) + 5x + 2 – 3x + 3y + 2 what is the anwer

OpenStudy (anonymous):

distribute the negative into the ( )--it's like multiplying by -1. Gives: -2x+3y+5x+2-3x+3y+2 combine like terms--(-2x+5x+3x)=6x etc. do you understand?

OpenStudy (anonymous):

(3y+3y)=6y 2+2=4 Result 6x+6y+4